I have 2 questions. absolute value [3x-2]-9<5. The second one is absolute value[4x-6]-2>=5.
Thanks.

Answer by CharlesG2(834)

I have 2 questions. absolute value [3x-2]-9<5. The second one is absolute value[4x-6]-2>=5.
1)
|3x - 2| - 9 < 5
|3x - 2| < 14
-14 < 3x - 2 < 14
-14 + 2 < 3x < 14 + 2
-12 < 3x < 16
-4 < x < 16/3
2)
|4x - 6| - 2 >= 5
|4x - 6| >= 7
4x - 6 <= -7 or 4x - 6 >= 7
4x <= -1 or 4x >= 13
x <= -1/4 or x >= 13/4